# Target search optimization by threshold resetting

### v2

@@ Abstract
Introduction. Resetting mediated target search has been a focal point of scientific interest in statistical physics and stochastic processes. A hallmark feature of resetting is that it can expedite the completion of complex processes by periodically reinstating a search process to a known or random state, thus uncovering new pathways, avoiding potential obstacles and in effect, enhancing the efficiency in locating desired targets evans_stochastic_2020; pal2024random. The advent of resetting based optimization techniques has truly invigorated much theoretical evans_diffusion_2011; kusmierz2014first; pal2016diffusion; reuveni_optimal_2016; pal_first_2017; chechkin2018random; belan2018restart; pal2019first; kumar2023universal; de2023resetting; pal2019landau; sar2023resetting; bonomo2021mitigating; de2022optimal; bhat2016stochastic; campos2015phase; pal_search_2020; biswas2024search; nagar2016diffusion, experimental tal2020experimental; besga2020optimal; faisant2021optimal; altshuler2024environmental; paramanick2024uncovering and computational jolakoski2023first; blumer2024combining; munoz2025learning research of first-passage phenomena.
@@
The aim of this work is to introduce target search optimization under a threshold resetting (TR) mechanism. In contrast to the prototypical scenario where resetting is imposed by an external time clock, TR is an event driven phenomenon in which resetting occurs when the system operates under “safety covenant” or broadly, meets certain “threshold rules”. Thresholds play a crucial role in governing transitions across numerous systems, acting as critical points that trigger significant changes. In neuroscience, integrate-and-fire neurons accumulate stimuli until their membrane potential crosses a threshold, leading to an action potential before resetting to their initial state burkitt2006review; bachar2012stochastic; gerstein1964random; taillefumier2013phase. Financial strategies like stop-loss and take-profit rely on preset thresholds to limit losses or secure gains kaufman2013trading; shiryaev2007optimal; zhang2001stock; miller1966model. In software engineering, circuit breakers operate similar to threshold-based reset mechanisms, preventing an application to access an unresponsive server nygard2018release; surendro2021circuit; montesi2018decorator. In physics, fiber bundle models describe how materials distribute stress until a rupture threshold is met, after which the load is redistributed among the remaining fibers pradhan2010failure; hansen2015fiber. Simple stroboscopic threshold mechanisms have also been employed to obtain sustained temporal regularity in chaotic systems with applications to laser modeling sinha2001using; sinha1998dynamics; sinha1999computing; bhowmick2014targeting.
@@
In the recent past, De Bruyne et al investigated spatial properties of a single diffusing particle under TR mechanism (referred as first-passage resetting there) disclosing dynamic nonstationarity of the process de2020optimization; de2021optimization. Yet, another unexplored aspect of the problem is the target search optimization – the study of which is at the heart of this letter. To put things in perspective, consider a first passage process that is conducted by a group of agents, setting off from a certain location and looking for a target until it is found redner2001; bray2013persistence; metzler_first-passage_2014; martinez2013optimizing; bhattacharya2014collective; falcon2019collective. The process is, however, restarted at some random time whenever one of the agents reaches a pre-assigned threshold. At that time all the searchers are reset simultaneously to the initial locations from which they renew their search. This collective resetting couples the degrees of freedom and makes the system strongly correlated biroli2023extreme; biroli2023critical; biroli2024exact. Such collective TR mechanism has natural justifications as earlier instances of experiments with fish school couzin2005effective; couzin2011uninformed; conradt2005consensus and swarming robots brambilla2013swarm; lerman2004review have demonstrated event-triggered collective transitions—where a local individual event (such as the detection of a fault or hazard by a single agent) prompts a coordinated adjustment in navigation strategy or “reset” across the population. Similarly, in individuals or groups of agents such as clonal raider ants (Ooceraea biroi), sensory thresholds shaped by factors like environment, and external stimuli influence decisions vital for survival and reproduction chase2025physics. Inspired by these connections, this letter aims to analyze the statistical signatures of the random completion time for this correlated multi-agent search process under threshold controlled resetting. To this end, we develop a unified framework for generic governing dynamics and show that TR can also be used as a potent mechanism to facilitate improved target search dynamics and efficiency. Finally, inspired from the control-management problem, we define a cost objective function that interpolates between the successful search time and the effective penalty owing to resetting events and examine whether this can be minimized.
@@
Since the searchers are non-interacting and identical in nature, we can write $q_{N}(L,x_{0},t)=\left[Q(L,x_{0},t)\right]^{N}$, where $Q(L,x_{0},t)$ is the survival probability of a single searcher, starting at $x_{0}\in[0,L]$ upto time $t$, with absorbing boundaries at both 0 and $L$ majumdar2020extreme. On the other hand, $j_{L,N}(L,x_{0},t)$ consists of current to the resetting threshold from a single searcher i.e. $j_{L,1}(L,x_{0},t)$ barring the rest N1N-1 of them which survive with probability $\left[Q(L,x_{0},t)\right]^{N-1}$ so that $j_{L,N}(L,x_{0},t)=Nj_{L,1}(L,x_{0},t)\left[Q(L,x_{0},t)\right]^{N-1}$ (the factor $N$ accounts for the fact that any of the $N$ searchers can contribute to $j_{L,1}(L,x_{0},t)$). Substituting these statistical metrics into Eq. (3), we arrive at
@@
Ballistic search under TR mechanism. To illustrate the theoretical framework developed above, we examine the paradigmatic case of ballistic searchers with randomized velocity. Such models have been commonly used to describe navigation pattern for some foragers such as microzooplankton, jackals and marine predators reynolds2009optimising; sims2008scaling, to scrutinize optimal strategies for identifying randomly located target sites santos2004optimal; james2008optimizing and also in statistical physics villarroel2018continuous; biroli2024exact; martens2012probability; sprenger2023dynamics; bechinger2016active. Crucially, its analytical tractability offers valuable insights that enhance our understanding of TR-facilitated target search processes. Subsequently, we have extended our analysis to diffusive searchers.
@@
In the limit $u\to 1$ when the resetting threshold is at the closest proximity to $x_{0}$, the resetting events are frequent and we find $\langle{\mathcal{N}}_{\text{TR}}(u\to 1,N)\rangle=2^{N}-1$ SI. However, since all such resetting events consume almost no time with $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u\to 1)\rangle\to 0$, the cumulative contribution to the MFPT from TR events becomes almost negligible. Notably, the finite contribution to the MFPT in this case comes solely from those correlated events where all the ballistic searchers collectively sample negative velocities right after the resetting or at the inception to drift right toward the target, marking a successful completion. Thus, $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u\to 1)\rangle$ effectively contributes to the MFPT whose asymptotic form is given by Eq. (10). The typical trajectories in this limit (see the bottom-right panel in Fig. 2b) confirm this observation. Quite interestingly, this emergent global optimality at $u\to 1$ aligns with a widely used threshold control strategy in chaos theory sinha2001using; bhowmick2014targeting.
@@
Figure 3: Behavior of the cost function with respect to $u=x_{0}/L$ for $N=3$ and $\beta=1$. The cost attains a minimum at an optimal $u^{*}$ – the variation of which as a function of NN is shown in the inset panel. The squares represent data from simulations.
Cost of TR mechanism. A central quest in the resetting discipline is to harness the cost, be dynamic de2020optimization; sunil2023cost; sunil2024minimizing or thermodynamic tal2020experimental; tal2025smart; olsen2024thermodynamic; gupta2020work; pal2023thermodynamic, of the overall process. Imitating the penalty during a power grid breakdown ajjarapu1994optimal; feng2000comprehensive, we posit, for each resetting event, a cost that is proportional to the MFPT as well as the total number of TR events. Henceforth, the goal is to determine the optimal operation of this system, for which the cost function
@@
Conclusions. Motivated by threshold-driven phenomena that refer to processes or events that occur only when a system surpasses a critical threshold, we study the first passage properties of an arbitrary search dynamics under threshold resetting mechanism. We develop a general theoretical framework and study the optimization properties of the MFPT which is a useful metric in learning the efficiency of a target search process. The optimization reported here also resonates with the performance maximization by the agents (robots or foraging animals) that collaborate and achieve tasks (for instance finding the desired target between many competitive ones by a fish school in couzin2011uninformed) in an optimal way.
In a classical foraging model where the searchers move ballistically, we discover a non-trivial optimization that suggests that persistent resetting to the threshold can render the MFPT globally minimum – an effect reminiscent to the threshold mechanism in nonlinear chaotic systems sinha2001using; bhowmick2014targeting. Extending the analysis to $N$ diffusive searchers also reveals rich optimization behaviors of the MFPT SI; diffusive_TR. By defining a cost function for such systems, we further demonstrate that the collective TR mechanism can achieve optimal cost.
To extend the formalism to capture more realistic scenarios, we can incorporate an overhead time for information transfer. This can be done in two ways: First, we can add a random (or deterministic) information passing time $\mathcal{T}^{O}$ whenever a searcher hits the threshold, after which all of them are reset to the initial position. In this case, one can show that the overall MFPT of the system would increase (however, this computation excludes the time required to transfer information to the others once a searcher successfully locates the target) diffusive_TR. A more involved extension would be to consider space–time coupled return protocols, where after a TR event searchers return to the initial configuration via finite-time spatial dynamics pal_search_2020; tal2020experimental; biswas2024search.
Finally, it is worth noting that the formalism can also be applied to extract higher order moments of the search time in higher dimensions and to the dynamics of non-coordinate variables like a fluctuating potential as in the case of firing neurons where the membrane potential of an individual synapse can reset after reaching a critical value burkitt2006review; taillefumier2013phase. Ongoing works also include characterizing nonequilibrium stationary states for such long-range systems under TR diffusive_TR; BMS2025. Given its simple implementation yet achieving target functionality makes the TR mechanism a promising avenue in the target search problem which can have far reaching implications beyond statistical physics to foraging, control theory, finance, structural failure-repair analysis and in computing systems.
@@
End Matter
Diffusive searchers under TR. Let us consider the case of $N$ diffusive searchers with the same set-up as for the ballistic walkers. To proceed, we will need the single searcher diffusive propagator $G(x,t|x_{0})$ in the presence of two absorbing boundaries at $x=0$ and $x=L$. This is well known in literature and is given by redner2001
$G(x,t|x_{0})=\frac{2}{L}\sum_{n=0}^{\infty}\sin\left(\frac{\pi nx}{L}\right)\sin\left(\frac{\pi nx_{0}}{L}\right)e^{-\frac{n^{2}\pi^{2}Dt}{L^{2}}},$ (13)
where $D$ is the diffusion constant. Using this, we can calculate the survival probability and the currents through the threshold and the target in terms of two quantities $u$ and $\tau_{d}=L^{2}/D$ so that SI
$Q(u,t)=\frac{2}{\pi}\sum_{n=1}^{\infty}\left(\frac{1-(-1)^{n}}{n}\right)\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (14)
$j_{0,1}(u,t)=\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (15)
$j_{L,1}(u,t)=-\frac{2\pi}{\tau_{d}}\sum_{n=1}^{\infty}n(-1)^{n}\sin\left(n\pi u\right)e^{-\frac{n^{2}\pi^{2}t}{\tau_{d}}},$ (16)
Figure 4: Diffusive searchers under TR: Mean first-passage time (MFPT) as a function of the threshold parameter $u$ for different values of NN as in Eq. (17). The horizontal dashed lines (for $N\geq 3$) represent the MFPT in the absence of threshold resetting, corresponding to the $u\to 0$ limit (note that the MFPTs for $N=1,2$ in this limit are diverging). The fact that the MFPT curves fall below these baselines highlights the enhanced search efficiency achieved through the TR mechanism, also in the case of diffusive searchers. The markers represent the simulation results.
The MFPT can be written in a dimensionless form as $\langle\overline{\mathcal{T}}_{N}^{\text{TR}}(u)\rangle=\langle\mathcal{T}_{N}^{\text{TR}}(x_{0},L)\rangle/\tau_{d}=\mathcal{F}(u,N)$. Skipping details from SI, we obtain an exact expression of the scaling function $\mathcal{F}(u,N)$ representing the MFPT which is given by

### v3

# Valid PACS appear here

@@ Abstract
Introduction. Resetting mediated target search has been a focal point of scientific interest in statistical physics and stochastic processes. A hallmark feature of resetting is that it can expedite the completion of complex processes by periodically reinstating a search process to a known or random state, thus uncovering new pathways, avoiding potential obstacles and in effect, enhancing the efficiency in locating desired targets Evans et al. (2020); Pal et al. (2024). The advent of resetting based optimization techniques has truly invigorated much theoretical Evans and Majumdar (2011); Kusmierz et al. (2014); Pal et al. (2016); Reuveni (2016); Pal and Reuveni (2017); Chechkin and Sokolov (2018); Belan (2018); Pal et al. (2019); Kumar and Pal (2023); De Bruyne and Mori (2023); Pal and Prasad (2019); Sar et al. (2023); Bonomo et al. (2022); De Bruyne et al. (2022); Bhat et al. (2016); Campos and Méndez (2015); Pal et al. (2020); Biswas et al. (2024); Nagar and Gupta (2016), experimental Tal-Friedman et al. (2020); Besga et al. (2020); Faisant et al. (2021); Altshuler et al. (2024); Paramanick et al. (2024) and computational Jolakoski et al. (2023); Blumer et al. (2024); Muñoz-Gil et al. (2025) research of first-passage phenomena.
@@
The aim of this work is to introduce target search optimization under a threshold resetting (TR) mechanism. In contrast to the prototypical scenario where resetting is imposed by an external time clock, TR is an event driven phenomenon in which resetting occurs when the system operates under “safety covenant” or broadly, meets certain “threshold rules”. Thresholds play a crucial role in governing transitions across numerous systems, acting as critical points that trigger significant changes. In neuroscience, integrate-and-fire neurons accumulate stimuli until their membrane potential crosses a threshold, leading to an action potential before resetting to their initial state Burkitt (2006); Bachar et al. (2012); Gerstein and Mandelbrot (1964); Taillefumier and Magnasco (2013). Financial strategies like stop-loss and take-profit rely on preset thresholds to limit losses or secure gains Kaufman (2013); Shiryaev (2007); Zhang (2001); Miller and Orr (1966). In software engineering, circuit breakers operate similar to threshold-based reset mechanisms, preventing an application to access an unresponsive server Nygard (2018); Surendro et al. (2021); Montesi and Weber (2018). In physics, fiber bundle models describe how materials distribute stress until a rupture threshold is met, after which the load is redistributed among the remaining fibers Pradhan et al. (2010); Hansen et al. (2015). Simple stroboscopic threshold mechanisms have also been employed to obtain sustained temporal regularity in chaotic systems with applications to laser modeling Sinha (2001); Sinha and Ditto (1998, 1999); Bhowmick et al. (2014).
@@
In the recent past, De Bruyne et al investigated spatial properties of a single diffusing particle under TR mechanism (referred as first-passage resetting there) disclosing dynamic nonstationarity of the process De Bruyne et al. (2020); De Bruyne and Redner (2021). Yet, another unexplored aspect of the problem is the target search optimization – the study of which is at the heart of this letter. To put things in perspective, consider a first passage process that is conducted by a group of agents, setting off from a certain location and looking for a target until it is found Redner (2001); Bray et al. (2013); Metzler et al. (2014); Martínez-García et al. (2013); Bhattacharya and Vicsek (2014); Falcón-Cortés et al. (2019). The process is, however, restarted at some random time whenever one of the agents reaches a pre-assigned threshold. At that time all the searchers are reset simultaneously to the initial locations from which they renew their search. This collective resetting couples the degrees of freedom and makes the system strongly correlated Biroli et al. (2023a, b, 2024). Such collective TR mechanism has natural justifications as earlier instances of experiments with fish school Couzin et al. (2005, 2011); Conradt and Roper (2005) and swarming robots Brambilla et al. (2013); Lerman et al. (2004) have demonstrated event-triggered collective transitions—where a local individual event (such as the detection of a fault or hazard by a single agent) prompts a coordinated adjustment in navigation strategy or “reset” across the population. Similarly, in individuals or groups of agents such as clonal raider ants (Ooceraea biroi), sensory thresholds shaped by factors like environment, and external stimuli influence decisions vital for survival and reproduction Chase and Peleg (2025). Inspired by these connections, this letter aims to analyze the statistical signatures of the random completion time for this correlated multi-agent search process under threshold controlled resetting. To this end, we develop a unified framework for generic governing dynamics and show that TR can also be used as a potent mechanism to facilitate improved target search dynamics and efficiency. Finally, inspired from the control-management problem, we define a cost objective function that interpolates between the successful search time and the effective penalty owing to resetting events and examine whether this can be minimized.
@@
Since the searchers are non-interacting and identical in nature, we can write $q_{N}(L,x_{0},t)=\left[Q(L,x_{0},t)\right]^{N}$, where $Q(L,x_{0},t)$ is the survival probability of a single searcher, starting at $x_{0}\in[0,L]$ upto time $t$, with absorbing boundaries at both 0 and $L$ Majumdar et al. (2020). On the other hand, $j_{L,N}(L,x_{0},t)$ consists of current to the resetting threshold from a single searcher i.e. $j_{L,1}(L,x_{0},t)$ barring the rest N1N-1 of them which survive with probability $\left[Q(L,x_{0},t)\right]^{N-1}$ so that $j_{L,N}(L,x_{0},t)=Nj_{L,1}(L,x_{0},t)\left[Q(L,x_{0},t)\right]^{N-1}$ (the factor $N$ accounts for the fact that any of the $N$ searchers can contribute to $j_{L,1}(L,x_{0},t)$). Substituting these statistical metrics into Eq. (3), we arrive at
@@
Ballistic search under TR mechanism. To illustrate the theoretical framework developed above, we examine the paradigmatic case of ballistic searchers with randomized velocity. Such models have been commonly used to describe navigation pattern for some foragers such as microzooplankton, jackals and marine predators Reynolds and Bartumeus (2009); Sims et al. (2008), to scrutinize optimal strategies for identifying randomly located target sites Santos et al. (2004); James et al. (2008) and also in statistical physics Villarroel and Montero (2018); Biroli et al. (2024); Martens et al. (2012); Sprenger et al. (2023); Bechinger et al. (2016). Crucially, its analytical tractability offers valuable insights that enhance our understanding of TR-facilitated target search processes. Subsequently, we have extended our analysis to diffusive searchers.
@@
In the limit $u\to 1$ when the resetting threshold is at the closest proximity to $x_{0}$, the resetting events are frequent and we find $\langle{\mathcal{N}}_{\text{TR}}(u\to 1,N)\rangle=2^{N}-1$ SI . However, since all such resetting events consume almost no time with $\langle\overline{\mathbb{T}}_{N}^{\text{TR}}(u\to 1)\rangle\to 0$, the cumulative contribution to the MFPT from TR events becomes almost negligible. Notably, the finite contribution to the MFPT in this case comes solely from those correlated events where all the ballistic searchers collectively sample negative velocities right after the resetting or at the inception to drift right toward the target, marking a successful completion. Thus, $\langle\overline{\mathbb{T}}_{N}^{\text{FP}}(u\to 1)\rangle$ effectively contributes to the MFPT whose asymptotic form is given by Eq. (10). The typical trajectories in this limit (see the bottom-right panel in Fig. 2b) confirm this observation. Quite interestingly, this emergent global optimality at $u\to 1$ aligns with a widely used threshold control strategy in chaos theory Sinha (2001); Bhowmick et al. (2014).
@@
Figure 3: Behavior of the cost function with respect to $u=x_{0}/L$ for $N=3$ and $\beta=1$. The cost attains a minimum at an optimal $u^{*}$ – the variation of which as a function of $N$ is shown in the inset panel. The squares represent data from simulations.
Cost of TR mechanism. A central quest in the resetting discipline is to harness the cost, be dynamic De Bruyne et al. (2020); Sunil et al. (2023, 2024) or thermodynamic Tal-Friedman et al. (2020, 2025); Olsen et al. (2024); Gupta et al. (2020); Pal et al. (2023), of the overall process. Imitating the penalty during a power grid breakdown Ajjarapu et al. (1994); Feng et al. (2000), we posit, for each resetting event, a cost that is proportional to the MFPT as well as the total number of TR events. Henceforth, the goal is to determine the optimal operation of this system, for which the cost function
@@
Conclusions. Motivated by threshold-driven phenomena that refer to processes or events that occur only when a system surpasses a critical threshold, we study the first passage properties of an arbitrary search dynamics under threshold resetting mechanism. We develop a general theoretical framework and study the optimization properties of the MFPT which is a useful metric in learning the efficiency of a target search process. The optimization reported here also resonates with the performance maximization by the agents (robots or foraging animals) that collaborate and achieve tasks (for instance finding the desired target between many competitive ones by a fish school in Couzin et al. (2011)) in an optimal way.
In a classical foraging model where the searchers move ballistically, we discover a non-trivial optimization that suggests that persistent resetting to the threshold can render the MFPT globally minimum – an effect reminiscent to the threshold mechanism in nonlinear chaotic systems Sinha (2001); Bhowmick et al. (2014). Extending the analysis to $N$ diffusive searchers also reveals rich optimization behaviors of the MFPT SI ; Biswas et al. (2025). By defining a cost function for such systems, we further demonstrate that the collective TR mechanism can achieve optimal cost.
To extend the formalism to capture more realistic scenarios, we can incorporate an overhead time for information transfer. This can be done in two ways: First, we can add a random (or deterministic) information passing time $\mathcal{T}^{O}$ whenever a searcher hits the threshold, after which all of them are reset to the initial position. In this case, one can show that the overall MFPT of the system would increase (however, this computation excludes the time required to transfer information to the others once a searcher successfully locates the target) Biswas et al. (2025). A more involved extension would be to consider space–time coupled return protocols, where after a TR event searchers return to the initial configuration via finite-time spatial dynamics Pal et al. (2020); Tal-Friedman et al. (2020); Biswas et al. (2024).
Finally, it is worth noting that the formalism can also be applied to extract higher order moments of the search time in higher dimensions and to the dynamics of non-coordinate variables like a fluctuating potential as in the case of firing neurons where the membrane potential of an individual synapse can reset after reaching a critical value Burkitt (2006); Taillefumier and Magnasco (2013). Ongoing works also include characterizing nonequilibrium stationary states for such long-range systems under TR Biswas et al. (2025); Biroli et al. (2025). Given its simple implementation yet achieving target functionality makes the TR mechanism a promising avenue in the target search problem which can have far reaching implications beyond statistical physics to foraging, control theory, finance, structural failure-repair analysis and in computing systems.
@@
Supplemental material for
“Target search optimization by threshold resetting”
This Supplemental Material provides detailed derivations of some of our main results highlighted in the Letter. Moreover, it also provides additional discussions that support our finding as announced in the Letter.
S1 Survival probability and probability current of a single ballistic searcher
In this section, we compute the survival probability $Q(L,x_{0},t)$ and the current through the resetting threshold $j_{L,1}(L,x_{0},t)$ for a single ballistic searcher as used in Eq. (5) and (6) of the main text. If $v_{i}$ denotes the random initial velocity of the $i^{th}$ searcher then it can survive up to time $t$ without hitting either target or resetting threshold in two ways: First, if $v_{i}t<L-x_{0}$ when $v_{i}>0$ so that the walker starts moving towards the resetting threshold but failing to reach the same in the given time and second, when $|v_{i}|t<x_{0}$ if the initial velocity $v_{i}<0$ so that the walker starts moving towards the target yet fails to hit it by time $t$. Summing over these two possibilities, we can write the survival probability as
$Q(L,x_{0},t)=\int_{0}^{\infty}dv~\phi(v)~\theta(L-x_{0}-vt)+\int_{-\infty}^{0}dv~\phi(v)~\theta(x_{0}-|v|t),$ (S1)
where $\theta(z)$ is the Heaviside theta function which takes values unity only when $z>0$, and 0 otherwise. Assuming $\phi(v)$ to be symmetric so that $\phi(-v)=\phi(v)$, one can rewrite the above equation as
$Q(L,x_{0},t)=\int_{0}^{\frac{L-x_{0}}{t}}dv~\phi(v)+\int_{0}^{\frac{x_{0}}{t}}dv~\phi(v).$ (S2)
Denoting the function $\Phi(v)$ as
$\Phi(v)=\int_{0}^{v}dv^{\prime}\phi(v^{\prime}),$ (S3)
we finally arrive at
$Q(L,x_{0},t)=\Phi\left(\frac{L-x_{0}}{t}\right)+\Phi\left(\frac{x_{0}}{t}\right),$ (S4)
which is Eq. (5) of the main text. The probability current through the resetting threshold at $x=L$ is essentially the total number of accumulated particles there at time $t$ so that
$j_{L,1}(L,x_{0},t)$ $=\int_{0}^{\infty}dv~\phi(v)~\delta\left(t-\frac{L-x_{0}}{v}\right)$
$=\frac{L-x_{0}}{t^{2}}~\phi\left(\frac{L-x_{0}}{t}\right),$ (S5)
which is Eq. (6) of the main text. Here, $\delta(x)$ is the Dirac-delta function. Note that only the positive velocities contribute to the current through the threshold while the negative velocities contribute to the current through the target at the origin given by
$j_{0,1}(L,x_{0},t)$ $=\int_{-\infty}^{0}dv~\phi(v)~\delta\left(t-\frac{x_{0}}{v}\right)$
$=\frac{x_{0}}{t^{2}}~\phi\left(\frac{x_{0}}{t}\right).$ (S6)
S2 MFPT of a single ballistic searcher under TR Mechanism Eq. (7)
This section extends the derivation of the MFPT of a single searcher under TR mechanism which was announced in Eq. (7) of the main text. To this end, we start by deriving the numerator in Eq. (3) of the main text by setting $N=1$ which yields the unconditional MFPT for a single particle to reach either the resetting threshold or the target
$\langle T_{1}(L,x_{0})\rangle$ $=\int_{0}^{\infty}dt~Q(L,x_{0},t)$
$=\int_{0}^{\infty}dt\int_{0}^{\infty}dv~\phi(v)~\theta(L-x_{0}-vt)+\int_{0}^{\infty}dt\int_{-\infty}^{0}dv~\phi(v)~\theta(x_{0}-|v|t)$
$=\int_{0}^{\infty}dv~\phi(v)\int_{0}^{\infty}dt~\theta(L-x_{0}-vt)+\int_{-\infty}^{0}dv~\phi(v)~\int_{0}^{\infty}dt\theta(x_{0}-|v|t)$
$=\int_{0}^{\infty}dv~\phi(v)~\left(\frac{L-x_{0}}{v}\right)+\int_{-\infty}^{0}dv~\phi(v)~\left(\frac{x_{0}}{|v|}\right).$ (S7)
Using the symmetry property of the velocity distribution i.e. $\phi(-v)=\phi(v)$, we can further simplify the above equation to find
$\langle T_{1}(L,x_{0})\rangle$ $=\int_{0}^{\infty}dv~\phi(v)~\left(\frac{L-x_{0}}{v}\right)+\int_{0}^{\infty}dv~\phi(v)~\left(\frac{x_{0}}{v}\right)$
$=L\int_{0}^{\infty}dv~\frac{\phi(v)}{v}.$ (S8)
Let us turn our attention to compute the denominator of Eq. (3) from the main text for $N=1$. We first note that it is just the splitting probability $\epsilon_{0}(L,x_{0},1)$ of a single searcher to reach the target located at the origin before hitting the threshold. For a symmetric velocity distribution, $\epsilon_{0}(L,x_{0},1)$ is simply $1/2$ since any searcher can acquire a positive (negative) velocity with probability $1/2$ indifferent to the initial coordinate $x_{0}$. This can also be seen formally as
$\epsilon_{0}(L,x_{0},1)$ $=1-\int_{0}^{\infty}dt~j_{L,1}(L,x_{0},t)$
$=\int_{0}^{\infty}dt~j_{0,1}(L,x_{0},t)$
$=\int_{0}^{\infty}dt~\frac{x_{0}}{t^{2}}~\phi(x_{0}/t)$
$=\frac{1}{2}.$ (S9)
Putting together both the numerator and denominator in Eq. (3) of the main text, the MFPT under TR mechanism for a single searcher takes the form
$\langle\mathcal{T}_{1}^{\text{TR}}(L,x_{0})\rangle=\frac{\langle T_{1}(L,x_{0})\rangle}{\epsilon_{0}(L,x_{0},1)}=2L\int_{0}^{\infty}dv~\frac{\phi(v)}{v},$ (S10)
which is Eq. (7) of the main text.
S3 Derivation of the scaling function ${\cal F}(u,N)$ as in Eq. (8)
In this section, we derive an exact form of the scaling function ${\cal F}(u,N)$ which was announced in the main text. To this end, we use Eq. (4) from the main text and note that we need the single particle survival probability (as in Eq. (S4)) and the probability current to the resetting threshold (as in Eq. (S5)) for the exponential velocity distribution $\phi(v)=\frac{1}{2v_{0}}e^{-|v|/v_{0}}$. In this case, we have $\Phi(v)=\frac{1}{2}\left(1-e^{-v/v_{0}}\right)$. Defining the following variables
$u=x_{0}/L,~~\tau_{b}=L/v_{0},$ (S11)
where $\tau_{b}$ is the timescale of a ballistic walker with speed $v_{0}$ and using Eq. (S4) and Eq. (S5), we can now compute the survival probability and the currents to the resetting threshold and the target respectively which take the following form
$Q(u,t)=1-\frac{1}{2}\exp\left({-\frac{\tau_{b}(1-u)}{t}}\right)-\frac{1}{2}\exp\left({-\frac{\tau_{b}u}{t}}\right),$ (S12)
$j_{L,1}(u,t)=\frac{\tau_{b}(1-u)}{2t^{2}}\exp\left({-\frac{\tau_{b}(1-u)}{t}}\right),$ (S13)
$j_{0,1}(u,t)=\frac{\tau_{b}u}{2t^{2}}\exp\left({-\frac{\tau_{b}u}{t}}\right).$ (S14)
We now recall from the main text that the splitting probability for any of the $N$ particles to hit the target is given by
$\epsilon_{0}(L,x_{0},N)$ $=1-\int_{0}^{\infty}dt~j_{L,N}(L,x_{0},t)$
$=1-\int_{0}^{\infty}dt~Nj_{L,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}$
$=\int_{0}^{\infty}dt~Nj_{0,1}(L,x_{0},t)[Q(L,x_{0},t)]^{N-1}.$ (S15)
Plugging the results from Eq. (S12)-(S15) into Eq. (4) of the main text, we find